OpenStudy (anonymous):

1.)A bag has 7 red marbles, 8 blue marbles and 3 black marbles. What is the probability of picking a blue marble, replacing it, and then picking a red marble? Answer a. 14/81 b. 7/108 c. 2/27 d. 1/9 2.)A bag has 7 red marbles, 8 blue marbles and 3 black marbles. What is the probability of picking a black marble, replacing it, and then picking a red marble? Answer a. 2/27 b. 7/108 c. 14/81 d. 1/9

OpenStudy (anonymous):

Simple formula: total of *color* / total

OpenStudy (kirbykirby):

\[\frac{8}{18}*\frac{7}{18}\]

OpenStudy (kirbykirby):

\(P(blue)=8/18\) since there are 18 marbles in total, and only 8 are blue Since you put back the marble, there are again 18 marbles in the bag. Now \(P(red)=7/18\) retricethere are 7 blue marbles out of 18 Now, you want prob(blue) AND prob(red), so you multiply the two probabilities by the product rule

OpenStudy (kirbykirby):

With that you should be able to do #2) since it's really the same problem, but just different colours are asked

OpenStudy (anonymous):

thank you!

OpenStudy (kirbykirby):

:)